OpenStudy (gamenerd123):

Simplify. √18 3√6 2√3 2√9 3√2

OpenStudy (londonlove2002):

wrong section

OpenStudy (gamenerd123):

??

Awolflover1 (awolflover1):

nope this is the right section haha

OpenStudy (peachpi):

do you know how to find the prime factorization of 18?

OpenStudy (londonlove2002):

its supposed to b in mathematics

OpenStudy (gamenerd123):

Well I'm actually trying to simplify it

OpenStudy (londonlove2002):

o sorry im in the wrong section

OpenStudy (jtm-pa):

\[3\sqrt{2}\] because 9*2 = 18 and 9s square root is 3 so you can remove that from the inside of the equation

OpenStudy (londonlove2002):

lol i feel stupid

OpenStudy (gamenerd123):

No need to Love :)

OpenStudy (gamenerd123):

But they are trying to find the perfect Square

OpenStudy (gamenerd123):

The perfect square that divides into 18

alones (alones):

wait so you're saying \(\ 3\sqrt{2}\) isn't an answer?

OpenStudy (samantha2234):

3 (sqrt)2

OpenStudy (gamenerd123):

Ok thx, can u help me with some more?